OpenStudy (anonymous):

A race pilot's average rate of speed over a 720-mile course is inversely proportional to the time in minutes t the pilot takes to fly a complete race course. The pilot's final score s is the average speed minus any penalty points p earned. 1. Write a function to model the pilot's score for a given t and p (Hint: d = rt) 2. Graph the function for a pilot who has two penalty points. 3. What is the maximum time a pilot with 2 penalty points can take to finish the course and still earn a score of at least 3?

OpenStudy (anonymous):

i really don't understand this, please help

ganeshie8 (ganeshie8):

For part 1, use the given hint. \[d=rt\] notice that we have \(d=720\) here, so \[720=rt \implies r = \dfrac{720}{t}\]

ganeshie8 (ganeshie8):

`The pilot's final score s is the average speed minus any penalty points p earned.` \[s = r - p\] plugin the value of \(r\) above

ganeshie8 (ganeshie8):

\[s=\dfrac{720}{t}-p\] thats the function for part1

OpenStudy (anonymous):

ok thank you, I actually already understood that part

OpenStudy (anonymous):

it's number 2 i need help with

ganeshie8 (ganeshie8):

good, for part2, simply plugin \(p=2\) and graph the function

OpenStudy (anonymous):

oh, yeah of course

OpenStudy (anonymous):

so for part 3 do I just look for the maximum point on the graph?

ganeshie8 (ganeshie8):

nope, for part3, simply plugin \(s=3\) in the function and solve \(t\)

ganeshie8 (ganeshie8):

\[3=\dfrac{720}{t}-2\] solve \(t\)

OpenStudy (anonymous):

t=144

OpenStudy (anonymous):

so that's the maximum time

ganeshie8 (ganeshie8):

Yes thats the maximum time acceptable anything greater than that will reduce the score

OpenStudy (anonymous):

ok, thanks again!

ganeshie8 (ganeshie8):

yw